OpenStudy (anonymous):

Which of the following are equivalent to 160%? 0.16 1.6 1

OpenStudy (anonymous):

1.6

OpenStudy (anonymous):

dangit one sec all the awnser didnt go up

OpenStudy (anonymous):

It's 1.6 100%

OpenStudy (anonymous):

100% positive

OpenStudy (texaschic101):

1.6

OpenStudy (texaschic101):

change the percent to a decimal and move it two spaces to the left

OpenStudy (anonymous):

1/3/6........5/8 those are c aand d

OpenStudy (anonymous):

Give texas a badge

OpenStudy (texaschic101):

160% = 1.6

OpenStudy (anonymous):

1/3/6........5/8 those are c aand d

OpenStudy (texaschic101):

what is 1/3/6 ???

OpenStudy (anonymous):

one and three sixth

OpenStudy (texaschic101):

1 3/6 = 9/6 = 1.5 (not equivalent) 5/8 = .625 (not equivalent)

OpenStudy (anonymous):

its a multiple choice quetion to so are u shure

OpenStudy (texaschic101):

the only thing equivalent to 160% in this problem is 1.6

OpenStudy (anonymous):

okkkkkkkk